OpenStudy (anonymous):

Prove that:- cosAsin(B-C) + cosBsin(C-A) + cosCsin(A-B) =0

OpenStudy (anonymous):

Help

hartnn (hartnn):

expand sin (x-y) = sin x cos y - cos x sin y all 3 sin terms

OpenStudy (anonymous):

oh wait

hartnn (hartnn):

if i am not wrong , everything will cancel out

OpenStudy (anonymous):

yup

hartnn (hartnn):

cool :)

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

hehe , i was doing cos c + cos d and all of that

OpenStudy (anonymous):

Infact you can do that..

OpenStudy (anonymous):

That is simpler to apply here.

OpenStudy (anonymous):

You must know that: \[2\cos(x)\sin(y) = \sin(x+y) - \sin(x-y)\]

OpenStudy (anonymous):

yeah i know :)

OpenStudy (anonymous):

yes yes we can apply that , also cool:)

OpenStudy (anonymous):

Multiply and divide by 2 first, and then apply the formula there, the terms will easily cancel out...

hartnn (hartnn):

always good to know alternate methods :)